OpenStudy (anonymous):

How many pounds of hamburger worth $1.05 per pound must be mixed with 60 pounds of hamburger worth $0.90 per pound to produce hamburger worth $1.00 per pound?

OpenStudy (anonymous):

38.5

OpenStudy (anonymous):

pounds

OpenStudy (anonymous):

that says its not right:(

OpenStudy (anonymous):

really? thats what i got :P

OpenStudy (amistre64):

h@1.05 = (1.00-60(.90))/(1.05-.90) if i see it right

OpenStudy (amistre64):

apparently i dont see it right :)

OpenStudy (amistre64):

h2 - h1 = 60 h2 - h1(1.05) = 60(.90) h1= (60(.90)-60)/(-1.05 + 1) = 120

OpenStudy (amistre64):

120(1.05) + 60(.90) = 180 ?? i get yes

OpenStudy (anonymous):

180 at 1.00 a pound :)